OpenStudy (anonymous):

An airplane flew 3 hours with a 60 mph head wind. The return trip with a tail wind of the same speed took 2 hours. Find the speed of the plane in still air. 380 mph 340 mph 320 mph 300 mph

OpenStudy (anonymous):

It would be 300 mph! So D

OpenStudy (anonymous):

speed = x + 60 mph Distance = d miles t = d / ( x + 60 ) = 2 d = 2x + 120 AGAINST wind -------------------- speed = x - 60 mph distance = d miles t = d / ( x - 60 ) = 3 d = 3x - 180 2x + 120 = 3x - 180 300 = x